7) (20 pts) Let X(t)-At be a random process, such that A is N(0, 1). (b) Find the auto-correlation function of X, Rx(t1, t2) E[X(ti)X(t2)
Answer #1

A)

E(X(t))=E(At)=tE(A) =t*0=0

B)

E(X(t1)*X(t2))=E(At1*At2))=t1*t2*E(A^2)=t1*t2*(0^2+1^2)

=t1*t2